# A Collaborative Content Moderation Framework for Toxicity Detection based on Conformalized Estimates of Annotation Disagreement

### v2

@@ 3.3 Performance Metrics
namely, the True Positive Rate (TPR) associated to the conformalized annotation disagreement estimate resulting from the auxiliary modeling task.
@@
[…]
@@
The metrics presented so far measure how well the framework performs in a CM context. However, we have also introduced additional metrics to support the analysis of each research question. In RQ1, where we analyze the general performance and calibration of the models, we use F1 score, logistic loss, Expected Calibration Error (ECE), and Adaptive Calibration Error (ACE) [36] for classification, and Mean Absolute Error (MAE) and Mean Squared Error (MSE) for regression. For RQ2, where we measure model uncertainty in classification and regression, the following metrics are used:
@@ 4.2 RQ2: Effects of Annotation Disagreement Integration on Uncertainty Quantification
Since proper model calibration is closely related to accurate uncertainty quantification (UQ) provided by CP methods, we hypothesize that multitask settings could improve model uncertainty estimation for both STL classification (CLASS) and regression (REG). To quantify improvements in the primary task (STL CLASS), we focus specifically on the MURE metric, which assesses how well the uncertainty method predicts cases where the model is incorrect. In addition to this metric, we also consider measures commonly used to evaluate uncertainty estimation methods, including marginal coverage, the correlation between uncertainty and annotation disagreement scores $r_{I,d}$ (which refers to the Pearson correlation between the estimated annotation disagreement based on the confidence interval and the true annotation disagreement), and the certainty F1 score. To compare uncertainty improvements, we use the same classification and multitask models described in Section 4.1, calibrated using the same validation/calibration dataset and evaluated on the same test set. Each model is calibrated with the three uncertainty estimation methods introduced in Section 3.2: LAC, Class Conditional LAC, and CRC.
@@ 4.3 RQ3: Benefits of an Auxiliary Task versus Independent Models in Collaborative Content Moderation
Previous studies have assessed the efficiency of moderators and moderation systems using review efficiency (MURE) [5], which considers only model performance while disregarding label data quality and text ambiguity. To address these factors, our moderation framework includes an auxiliary task designed to measure this concept. Since this is a novel approach, we introduce a new metric, CARE, to account for the system’s capacity to detect ambiguous examples, as explained in Section 3.3. To determine if a comment is ambiguous, a threshold $\gamma$ must be selected that aligns with the desired maximum level of annotation ambiguity. Given that the quantity of uncertain examples varies depending on each model’s capabilities, which in turn affects the CARE metric, we used the False Positive Rate (FPR) at a given TPR/MURE value, a metric commonly used in other areas including Out-of-Distribution (OoD) detection [38]. Different TPR values are applied for RAC models due to the interval size, which affects the quality of the ambiguity estimation.
@@
  • STL CLASS (upper section of Figure 1): In this case, classification model’s uncertainty is used to measure how well the model performs in detecting ambiguous terms, as uncertainty is sometimes associated with hard examples—instances that the model has not learned well. Since the model is not calibrated for the task, we could not compare it under the same true positive rate (TPR) because it is only able to achieve a maximum TPR of around 45%.
@@ 5 Results and Discussion
In this section we present and analyse the results for each research question. In Section 5.1 we discuss our findings for RQ1, highlighting the benefits of our approach in terms of model calibration and performance. Section 5.2 focuses on the advantages of incorporating the auxiliary task for improving uncertainty estimation in the primary classification task. In Section 5.3 we compare the multitask and composite approaches for content moderation, with a focus on comment annotation disagreement. For each RQ, we begin by summarizing the key performance metrics and findings.
@@ 5.2 RQ2: Effects of Annotation Disagreement Integration on Uncertainty Quantification
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$uparrow
CLASS 0.4076 0.8046 0.901 0.2767
@@
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$uparrow
@@
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$uparrow
@@
Model name Method CP Interval size $\downarrow$ DI $\downarrow$ ICP $r_{I,d}$uparrow
@@ 5.3 RQ3: Benefits of an Auxiliary Task Versus Independent Models in Collaborative Content Moderation
REG Method Class CP Regre CP R-F1 (MTL/STL)$\uparrow$ FPR (MTL/STL)$\downarrow$ $r_{pb}$ (MTL/STL) $\uparrow$
@@ 6 Conclusions and Future Work
In this work, we introduced a novel framework for collaborative content moderation that accounts for model uncertainty and addresses annotation disagreement through the introduction of an auxiliary task. This approach excels at capturing the inherent subjectivity of toxic comments, providing a moderation scheme that reflects this intrinsic complexity. Additionally, we proposed two new metrics to evaluate the comment review process: F1 Review Efficiency and CARE. These metrics incorporate the concept of annotation disagreement into the review process, which the SOTA MURE does not consider.
Throughout the manuscript we have analysed the impact of incorporating this auxiliary task in terms of calibration, model performance, and uncertainty estimation for both the primary task (classification) and the auxiliary task (regression), using various regression losses and uncertainty quantification techniques. We have found out that the integration of the auxiliary task improves not only the classification task, but also enhances the auxiliary task in terms of calibration and performance.
By demonstrating improvements in uncertainty estimation through the use of annotation disagreement, we have opened up an interesting new area of research, where future studies could build on towards enhancing existing classification-based tasks with better uncertainty estimation techniques. Finally, we have evaluated the quantitative benefits of employing a multitask architecture over a single-task approach, obtaining promising results. The multitask approach outperformed the single-task model in detecting incorrect predictions, and also quantified disagreement in comments as specified by the chosen $\gamma$, namely, the targeted ambiguity threshold.
@@
  • RQ1: We have observed that, consistently with previous studies, the incorporation of an auxiliary task focused on annotation disagreement enhances the calibration and performance of the toxicity classification task. In the case of the improvements of performance in the auxiliary task, an overall improvement was observed for the MSE and MAE metrics.
  • RQ2: In terms of improving uncertainty estimation by adding an auxiliary task, this addition has been shown to enhance the classification model’s uncertainty estimation compared to a single-task classifier. Furthermore, we observed a better point-biserial correlation between the model’s uncertainty and annotation disagreement, indicating a stronger understanding of borderline cases. For the regression task (disagreement estimation), the confidence intervals generated were wider than those of the single-task regression model used for the same task, particularly for the RAC and BCE models. However, the MSE model did not demonstrate consistent performance differences, making the overall enhancement inconclusive and highly dependent on the appropriate configuration based on the selection of the uncertainty method and the regression loss.
  • RQ3: The multitask architecture generally outperforms CoM, which consist of a single-task classification model and a single-task regression model. Although in some cases the improvement reported in our results is small, the MTL framework is more computationally efficient and offered advantages regarding the responses to RQ1 and RQ2. Compared to the single-task classifier, the multitask model provides more flexibility for moderation, as it directly computes annotation ambiguity. However, we acknowledge that the uncertainty estimated for the single-task classifier is still reasonably aligned with comment ambiguity.
@@
Toxicity classification, and by extension content moderation, faces a significant challenge due to the inherent subjectivity of the task. This subjectivity complicates dataset creation and the annotation process, as we must account for both explicit and implicit (contextualized) toxicity. Since a few datasets provide prior comments or context for toxic comments, it becomes difficult to accurately assess the toxicity during annotation. This can lead to several issues in the annotation process, as outlined by Zhang et al. [23]. The main challenges include: rater heterogeneity, individual differences, variations in working patterns, difficulties in understanding comment complexity, unclear task descriptions, and issues related to randomness. These challenges directly impact our proposed moderation framework, which leverages annotation disagreement as a core element. Another limitation is that a few datasets guarantee a sufficient number of annotators, which restricts our approach due to economic constraints, as increasing the number of annotators significantly raises the costs derived from the annotation process. Finally, our approach may be limited by the heterogeneity and distribution of annotation disagreement, as datasets like Wiki Detox predominantly contain low-disagreement comments. This makes it more challenging to effectively model annotation disagreement and incorporate it to frameworks like the one proposed in this manuscript.
@@
As part of our future work, we intend to explore additional uncertainty estimation techniques to investigate whether the performance gains from introducing the auxiliary task extend to other uncertainty estimation methods. This will help in assessing the robustness and generalizability of our framework when incorporating other uncertainty estimation techniques. Furthermore, we aim to evaluate its applicability to other tasks characterized by subjective annotations, such as sentiment analysis, where annotation disagreement and ambiguity play a critical role. Additionally, we envision that our approach holds potential in the crossroads of active learning and generative text modelling, as it can generate more diverse datasets and eventually, improve the detection of evolving patterns in toxic language expression. In this envisioned approach, we will investigate how to effectively avoid known issues with synthetic data generation in fine-tuning loops, including the model degradation due to content autophagy [40].

### v3

@@ 3.3 Performance Metrics
namely, the True Positive Rate (TPR) associated to the conformalized annotation disagreement estimate resulting from the auxiliary modelling task.
@@
$\text{R-F1}=2\cdot\frac{\textup{MURE}\cdot\textup{CARE}}{\textup{MURE}+\textup{CARE}}.$ (13)
@@
The metrics presented so far measure how well the framework performs in a CM context. However, we have also introduced additional metrics to support the analysis of each research question. In RQ1, where we analyse the general performance and calibration of the models, we use F1 score, logistic loss, Expected Calibration Error (ECE), and Adaptive Calibration Error (ACE) [59] for classification, and Mean Absolute Error (MAE) and Mean Squared Error (MSE) for regression. For RQ2, where we measure model uncertainty in classification and regression, the following metrics are used:
@@ 4.2 RQ2: Effects of Annotation Disagreement Integration on Uncertainty Quantification
Since proper model calibration is closely related to accurate UQ provided by CP methods, we hypothesize that multitask settings could improve model uncertainty estimation for both STL classification (CLASS) and regression (REG). To quantify improvements in the primary task (STL CLASS), we focus specifically on the MURE metric, which assesses how well the uncertainty method predicts cases where the model is incorrect. In addition to this metric, we also consider measures commonly used to evaluate uncertainty estimation methods, including marginal coverage, the correlation between uncertainty and annotation disagreement scores $r_{I,d}$ (which refers to the Pearson correlation between the estimated annotation disagreement based on the confidence interval and the true annotation disagreement), and the certainty F1 score. To compare uncertainty improvements, we use the same classification and multitask models described in Section 4.1, calibrated using the same validation/calibration dataset and evaluated on the same test set. Each model is calibrated with the three uncertainty estimation methods introduced in Section 3.2: LAC, Class Conditional LAC, and CRC.
@@ 4.3 RQ3: Benefits of an Auxiliary Task versus Independent Models in Collaborative Content Moderation
Previous studies have assessed the efficiency of moderators and moderation systems using review efficiency (MURE) [5], which considers only model performance while disregarding label data quality and text ambiguity. To address these factors, our moderation framework includes an auxiliary task designed to measure this concept. Since this is a novel approach, we introduce a new metric, CARE, to account for the system’s capacity to detect ambiguous examples, as explained in Section 3.3. To determine if a comment is ambiguous, a threshold $\gamma$ must be selected that aligns with the desired maximum level of annotation ambiguity. Given that the quantity of uncertain examples varies depending on each model’s capabilities, which in turn affects the CARE metric, we used the False Positive Rate (FPR) at a given TPR/MURE value, a metric commonly used in other areas including Out-of-Distribution (OoD) detection [61]. For BCE and MSE, we apply a TPR of 85%, while for RAC, we use a TPR of 95%. Different TPR values are applied for RAC models due to the interval size, which affects the quality of the ambiguity estimation. These values were selected to provide a fair and consistent basis for comparing the performance of each approach. By choosing TPR values that balance precision and recall for each model, we ensure that we can effectively measure and compare each method’s ability to handle uncertainty across different models.
@@
  • STL CLASS (upper section of Figure 1): In this case, neural network’s uncertainty is used to measure how well the network performs in detecting ambiguous terms, as uncertainty is sometimes associated with hard examples—instances that the model has not learned well. Since the model is not calibrated for the task, we could not compare it under the same true positive rate (TPR) because it is only able to achieve a maximum TPR of around 45%.
@@ 5 Results and Discussion
In this section, we present and analyse the results for each research question. In Section 5.1 we discuss our findings for RQ1, highlighting the benefits of our approach in terms of model calibration and performance. Section 5.2 focuses on the advantages of incorporating the auxiliary task for improving uncertainty estimation in the primary classification task. In Section 5.3 we compare the multitask and composite approaches for content moderation, with a focus on comment annotation disagreement. For each RQ, we begin by summarizing the key performance metrics and findings.
@@ 5.2 RQ2: Effects of Annotation Disagreement Integration on Uncertainty Quantification
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$ $\uparrow$
STL CLASS 0.4076 0.8046 0.901 0.2767
@@
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$ $\uparrow$
@@
Model name MURE $\uparrow$ F1 score $\uparrow$ Marginal Coverage $r_{pb}$ $\uparrow$
@@
Model name Method CP Interval size $\downarrow$ DI $\downarrow$ ICP $r_{I,d}$ $\uparrow$
@@ 5.3 RQ3: Benefits of an Auxiliary Task Versus Independent Models in Collaborative Content Moderation
[…]
@@ 6 Conclusions and Future Work
In this work, we introduce a novel framework for collaborative content moderation that addresses annotation disagreement and accounts for model uncertainty using a multitask neural network architecture, with the latter included as an auxiliary task. This approach excels at capturing the inherent subjectivity of toxic comments, providing a moderation scheme that reflects this intrinsic complexity. Additionally, we proposed two new metrics to evaluate the comment review process: F1 Review Efficiency and CARE. These metrics incorporate the concept of annotation disagreement into the review process, which the state-of-the-art MURE score does not consider.
Throughout the manuscript we have analysed the impact of incorporating this auxiliary task into the multitask neural architecture in terms of calibration, model performance, and uncertainty estimation for both the primary task (classification) and the auxiliary task (regression), using various regression losses and uncertainty quantification techniques. We have found that the integration of the auxiliary task improves not only the classification task, but also enhances the auxiliary task in terms of calibration and performance.
By demonstrating improvements in uncertainty estimation through the use of annotation disagreement, we have opened up an interesting new area of research, where future studies could build on towards enhancing existing classification-based tasks with better uncertainty estimation techniques. Finally, we have evaluated the quantitative benefits of employing a multitask learning architecture over a single-task neural network, obtaining promising results. The multitask approach outperformed the single-task model in detecting incorrect predictions, and also quantified disagreement in comments as specified by the chosen $\gamma$, namely, the targeted ambiguity threshold.
@@
  • RQ1: We have observed that, consistently with previous studies, the incorporation of an auxiliary task focused on annotation disagreement enhances the calibration and performance of the toxicity classification task. While it is true that overall calibration and performance improved with RAC, the gain observed was limited. In the case of the improvements of performance in the auxiliary task, an overall improvement was observed for the MSE and MAE metrics.
  • RQ2: In terms of improving uncertainty estimation by adding an auxiliary task, this addition has been shown to enhance the classifier’s uncertainty estimation compared to a single-task neural network. Furthermore, we observed a better point-biserial correlation between the model’s uncertainty and annotation disagreement, indicating a stronger understanding of borderline cases. For the regression task (disagreement estimation), the confidence intervals generated were wider than those of the single-task regression model used for the same task, particularly for the RAC and BCE models. However, the MSE model did not demonstrate consistent performance differences, making the overall enhancement inconclusive and highly dependent on the appropriate configuration based on the selection of the uncertainty method and the regression loss.
  • RQ3: The multitask architecture generally outperforms CoM, which consist of two separate, single-task neural networks for classification and regression. Although in some cases the improvement reported in our results is small, the MTL framework is more computationally efficient and offered advantages regarding the responses to RQ1 and RQ2. Compared to the single-task neural network, the multitask architecture provides more flexibility for moderation, as it directly computes annotation ambiguity. However, we acknowledge that the uncertainty estimated for the single-task classifier is still reasonably aligned with comment ambiguity.
@@
Toxicity classification, and by extension content moderation, faces a significant challenge due to the inherent subjectivity of the task. This subjectivity complicates dataset creation and the annotation process, as we must account for both explicit and implicit (contextualized) toxicity. Since a few datasets provide prior comments or context for toxic comments, it becomes difficult to accurately assess the toxicity during annotation. This can lead to several issues in the annotation process, as outlined by Zhang et al. [26]. The main challenges include: rater heterogeneity, individual differences, variations in working patterns, difficulties in understanding comment complexity, unclear task descriptions, and issues related to randomness. These challenges directly impact our proposed moderation framework, which leverages annotation disagreement as a core element. Another limitation arises from the availability of high-quality annotations. Many datasets do not ensure a sufficient number of annotators per instance, often due to economic constraints, as increasing the number of annotators significantly raises annotation costs. Since our framework depends on capturing disagreement effectively, datasets with limited annotations may not provide enough perspectives to reveal meaningful variations in subjective judgments. Additionally, our use of a balanced toxic dataset means that we have not validated how our framework performs in other settings, such as imbalanced distributions of toxic and non-toxic comments, which are more representative of real-world moderation tasks. A key limitation lies in the distribution of annotation disagreements within existing datasets. Many datasets, like Wiki Detox, are dominated by comments with minimal disagreement, meaning that annotators rarely differ significantly in their assessments. However, our approach is designed to leverage situations where there is a high level of disagreement between annotators. Therefore, the full potential of our framework is best realized when applied to datasets that capture a broader range of annotator perspectives. That being said, even in cases with little to no disagreement, our method still functions effectively as a traditional toxicity detector learned from instances with majority-voted labels.
@@
As part of our future work, we intend to explore additional uncertainty estimation techniques to investigate whether the performance gains from introducing the auxiliary task extend to other methods, such as Evidential Learning, Epistemic Neural Networks, and Bayesian Deep Learning. This will help assess the robustness and generalizability of our framework when incorporating uncertainty estimation techniques suitable for deep neural networks. Furthermore, we aim to evaluate its applicability to other tasks characterized by subjective annotations, such as sentiment analysis, where annotation disagreement and ambiguity play a critical role. Moreover, we have observed that the slight class imbalance has been counteracted by the use of focal loss. In the case of a higher degree of imbalance, a potential avenue for future work could involve incorporating other techniques to address data imbalance, such as oversampling the minority class, undersampling the majority class, and using cost-sensitive loss functions, among others. Additionally, given the broad applicability of our proposal, we plan to further test it in real-world environments and are eager to collaborate with industry leaders. This collaboration will help us refine and adapt our approach based on practical feedback, ensuring its effectiveness and relevance in real-world moderation workflows. Finally, we envision that our approach holds potential in the crossroads of active learning and generative text modelling, as it can generate more diverse datasets and eventually, improve the detection of evolving patterns in toxic language expression. In this envisioned approach, we will investigate how to effectively avoid known issues with synthetic data generation in fine-tuning loops, including the model degradation due to content autophagy [62].